OpenStudy (anonymous):

Find the diagonal of a square whose sides are of the given measure. Given = 16" on this question i have to drag and drop numbers and its already giving me a 2 on the answer

OpenStudy (anonymous):

16^2+16^2=x^2

OpenStudy (anonymous):

x^2=512 \[x=\sqrt{256*2}=16\sqrt{2}\]

OpenStudy (anonymous):

thankyou=)